OpenStudy (anonymous):

Please help... A photograph is 4 in. longer than it is wide. If the frame is 2 in. wide, the combined area of the photograph and the frame is 252 in^2. Find the dimensions of the photograph without the frame.

OpenStudy (anonymous):

You can use a variable for wide or its length. like this wide = x length = x+4 so the area = x(x+4) 252= x^2 +4x so you can now easily solve it and find x 0 = x^2+4x -252